OpenStudy (ineedhelp10):

Find the inverse: f(x) = x^2

OpenStudy (ineedhelp10):

So i need the inverse function of f(x)=x^2 . I need help understanding it

OpenStudy (anonymous):

do you know what inverse means??

OpenStudy (ineedhelp10):

Yeah, its the opposite

OpenStudy (anonymous):

so do you know how the graph f(x) = x^2 looks like?

OpenStudy (ineedhelp10):

Not really

OpenStudy (anonymous):

 it looks like that

OpenStudy (ineedhelp10):

Oh ok, so how will that help me finding the inverse function? Like what are the steps to finding it?

OpenStudy (anonymous):

you just need to use the function \[\sqrt{y}\]

OpenStudy (ineedhelp10):

How will that help ?

OpenStudy (anonymous):

well that is the inverse equation so when you plug in the number the you squared into the squre root it will give you the inverse

OpenStudy (anonymous):

EX: Square:(-2)2 = 4 Inverse (Square Root): √(4) = 2